OpenStudy (anonymous):

Maria received a gift of cash from her grandmother. She put 2/3 of the money in her savings account, spent $25 on a pair of jeans, and spent half the remaining money on a DVD. If Maria had $18 left over, how much money did she receive from her grandmother?

OpenStudy (anonymous):

You didn't include the amount of money Maria had left over. I can't help you until I know that.

OpenStudy (anonymous):

lol i had to go back n retype it,mii badd

OpenStudy (anonymous):

18 + 18 = 36 36 + 25 = 61 61 * 3 = $183

OpenStudy (anonymous):

thanks dude